40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4623     Entry time: Wed May 4 13:45:56 2011
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: locking last night 

Last night I was trying to calibrate the MICH error signal and the actuators on BS and ITMs.

However I gave up taking the data because the MC locking was unstable. MC3 drifted a lot.

ELOG V3.1.3-